OpenStudy (anonymous):

What is the solution to the system of equations? 5x+4y=8 x-2y=10

OpenStudy (mehek.15):

solve for x in the second equation first so we can plug into first one

OpenStudy (photon336):

have you heard of substitution?

OpenStudy (anonymous):

no @Photon336

OpenStudy (photon336):

@michelleball477 you can solve for one variable and then plug it back into the second equation. I'll show you how based on what @Mehek.15 said

OpenStudy (photon336):

Let's solve for x in the second equation: x-2y = 10 to do this we need to get x by itself. we do this by adding 2y to both sides. x-2y = 10 This gives us x = 10 + 2y make sense so far?

OpenStudy (anonymous):

yea

OpenStudy (photon336):

omg. my equation bar froze..

OpenStudy (photon336):

what do you think we need to do now? x = 2y+10 5x+4y = 8

OpenStudy (anonymous):

idk let me work it out

OpenStudy (photon336):

@michelleball477 take a look at this we need to plug in 2y+10 wherever we see x so we plug this into 5x 5(2y+10)+4y = 8 now we distribute and get 10y+50+4y = 8 14y+50 = 8 14y = -42 y = -3

OpenStudy (photon336):

so now we have y = -3 we can put this into the second equation try to follow along here. now that we found y we can easily plug it back into the second equation to find x. x-2y = 10 x-2(-3) = 10 x+6 = 10 x = 4 so x= 4 and y = -3 and just to check let's plug this back into the top equation 5x+4y = 8 5(4)+4(-3) 50-12 = 8 so we know that this is our answer

OpenStudy (anonymous):

thanks i get it

OpenStudy (photon336):

try to look at this and work backwards, why do you think I substituted for one of the variables first?

OpenStudy (photon336):

@michelleball477